# Yixin Liu1 2  Lie Lu2  Jihui Jin2  Lichao Sun1  Andrea Fanelli2

### v2

@@ 5.1 Experimental Setup
Following prior works (San Roman et al., 2024; Chen et al., 2023), we use a sampling rate of 16 kHz and one-second mono samples for training ($T=16000$) under 16 diverse audio editing transformations. We train the models on a mixed audio dataset of 4100 hours containing speech (3016 hours VoxPopuli (Wang et al., 2021) and 100 hours LibriSpeech (Panayotov et al., 2015)), music (9 hours MusicCap (Agostinelli et al., 2023)) and sound effects (98 hours AudioSet (Gemmeke et al., 2017)). For evaluation, we use a held-out test set from MusicCap of size 100. The audio duration is set as 5s by default in evaluation. The loss weights are set as: $\lambda_{\text{TF}}=1$, $\lambda_{\text{adv}}=1$, $\lambda_{\ell_{1}}=0.1$, $\lambda_{\text{msspec}}=2$, $\lambda_{\text{detect}}=\lambda_{\text{message}}=10$. We use the Adam optimizer (Kingma & Ba, 2014) with learning rate 1e-5, $\beta_{1}=0.4$, $\beta_{2}=0.9$, and Exponential Moving Average (EMA) with decay factor of $0.99$ updated every step. The training uses batch size 16 for a total of 73k steps. The model’s latent dimension is set to $H=32$. For the attribution experiment, we follow the simulation protocol in San Roman et al. (2024) that defines a message pool of size $N$ ($\in\{100,1000,10000\}$), where each message is uniquely associated to a different user. As $N$ increases, the message length (in bits) also grows, making attribution more challenging due to the increased complexity of distinguishing individual messages. During decoding, the message is retrieved from the pool by selecting the one with the closest Hamming distance. We compare to four SoTA baselines: AudiowMark (Westerfeld, 2020), WavMark (Chen et al., 2023), TimbreWM (Liu et al., 2023a), AudioSeal (San Roman et al., 2024). Please refer to App. A for more details.
@@ 5.2 Detection and Attribution Effectiveness
Robustness to Standard Edits. We present the detection and attribution performance across a comprehensive suite of standard audio editing operations in Table 1. For detection, we report both the overall accuracy and the true/false positive rates (TPR/FPR). For attribution, we report the averaged performance across different user numbers (see Figure. 4 for decomposed result). Our method achieves new state-of-the-art performance on both tasks, maintaining high detection accuracy (99.19% average) and attribution accuracy (93%) on average across transformations. Notably, while AudioSeal demonstrates strong detection performance (97.1% average), it fails to perform effective attribution (39% average). The traditional approach AudiowMark exhibits more balanced detection-attribution trade-offs (around 88% attribution) but lower overall detection performance (92.9% average). On the speeding operation, our method achieves 99.5% detection accuracy while all other methods degrade to near random-guess levels (50-61%). Moreover, while all the models currently fail to attribute against speeding transformation, we demonstrate that our model can be further enhanced with a simple speed reversion layer to boost speed robustness while maintaining efficiency in the App. C.3.1. Finally, we validate the scalability of our watermarking system in Figure 4, which highlights that XAttnMark can maintain high attribution accuracy even when the users pool becomes larger. In summary, we demonstrate that XAttnMark is more robust under standard edits for both detection and attribution. See App. C.3 for more details.
@@
Robustness to Generative Edits. Beyond the standard audio transformations that are seen during training, audio generative editing is one particularly challenging type of editing that watermarking systems might have to endure at deployment (Liu et al., 2024b). To simulate this, we use two state-of-the-art audio generative models, AudioLDM2 (Liu et al., 2024b) and Stable Audio (Evans et al., 2024), with a text-guided DDIM inversion method proposed in ZETA (Manor & Michaeli, 2024). We test various editing strength $t\in\{10,70,110\}$, which represents the diffusion forward step when using DDIM inversion. As shown in Table 2, AudiowMark and WavMark degrade to a random-guess level performance, while TimbreWM and AudioSeal show inferior performance across different editing strengths around 50-60%. In contrast, our method maintains consistently high detection accuracy, averaging 91-94% across all editing strengths, and consistent across both generative models. This demonstrates that XAttnMark generalize better to unseen generative edits compared to existing methods. Please refer to the App. C.4 for more details.
Table 3: Detection performance and audio quality metrics after HSJA-based adversarial attacks on the waveform and spectrogram domains on Audiomark dataset (Liu et al., 2024a).
@@
Robustness to Adversarial Watermark Removal. As highlighted in Liu et al. (2024a), adversarial perturbations can be crafted to make the watermark undetectable, even with only black-box access to the detection model. To test robustness against these attacks, we employ the black-box HopSkipJumpAttack (HSJA) (Chen et al., 2020), to iteratively craft perturbations in the waveform or spectrogram domains that can fool the detector while minimizing the perturbation residual norm. As shown in Table 3, we first evaluate robustness against HSJA attacks in the two domains with query budget $Q=100$. The results show that, although the attack effectively decreases the detection accuracy, the quality of the perturbed audio is significantly degraded. Furthermore, with the waveform-based attack, increasing the query budget $Q$ from $100$ to $500$ further decreases the detection accuracy from 0.68 to 0.47. However, across different domains and query budgets, XAttnMark is consistently more robust than AudioSeal, maintaining higher detection accuracy and perceptual quality after the attack.
5.3 Quality and Stealthiness Assessment
Following San Roman et al. (2024), we evaluate the following objective quality metrics of the watermarked audio: Scale Invariant Signal to Noise Ratio (SI-SNR), as well as PESQ Rix et al. (2001), ViSQOL Hines et al. (2012) and STOI Taal et al. (2010). Furthermore, we also report the loudness of watermark residual $L(\bm{\delta_{w}})$ according to ITU-R BS.1770-4 (Series, 2011), which is an important metric for watermarking stealthiness, especially under residual-based detection (Yang et al., 2024). As shown in Table 4, XAttnMark achieves competitive performance across all perceptual quality metrics. While WavMark achieves the highest SI-SNR (36.35 dB), XAttnMark excels in perceptual quality measures, achieving the best STOI score (1.000) and the lowest watermark residual loudness (-54.63 dB). For the PESQ and ViSQOL scores, our method achieves competitive performance, with a PESQ score of 4.43 and a ViSQOL score of 4.56. Additionally, we conducted a subjective listening test of MUSHRA (Series, 2014) with 12 participants, and the results suggest that our method achieves a comparable quality with AudioSeal scoring around 91 (while ground truth scores around 95, please refer to the App. C.1 for more details). In summary, our method guarantees superior detection and attribution performance, while maintaining competitive perceptual quality. Please refer to the App. C.7 and App. C.8 for more discussion.
Table 4: Audio Quality Metrics. Comparison of perceptual quality metrics across different watermarking methods.
Methods SI-SNR $\uparrow$ PESQ $\uparrow$ STOI $\uparrow$ ViSQOL $\uparrow$ L($\delta_{x}$) $\downarrow$
@@ 5.4 Ablation Study
We first conduct ablation on the cross-attention module and temporal modulation layer to showcase their effectiveness for boosting learning in message decoding. As shown in Figure 4, our approach demonstrates superior learning efficiency, reaching approximately 98% accuracy. When removing the modulation component, performance drops significantly to around 60%. More dramatically, ablating the cross-attention mechanism causes accuracy to drop to random guess levels (50%), indicating this is a crucial architectural element. These results clearly demonstrate that while both components contribute to model performance, cross-attention is particularly vital for effective message retrieval learning. Furthermore, we ablate the proposed temporal-frequency loss $\mathcal{L}_{\text{TF}}$ and the adaptive bandwidth with constant weight $\gamma=1$. As shown in Table 4, both of them contribute to improving perceptual quality.
@@ 6 Conclusion
In this paper, we presented a neural audio watermarking framework that jointly achieves reliable watermark detection and accurate message attribution under challenging transformations, including generative editing. By integrating partial parameter sharing, a cross-attention detector, and a temporal conditioning module, we improved the robustness and decoding accuracy compared to prior neural watermarking methods. Our psychoacoustic-aligned time-frequency masking loss further enables quality-preserving watermarking. Though limitations remain for extreme transformations such as speeding, future work incorporating stronger temporal alignment could further improve real-world applicability.
@@
Our work advances robust audio watermarking to address critical challenges in media authenticity and intellectual property protection arising from generative AI. By enabling reliable detection and attribution of imperceptible watermarks across diverse transformations—including state-of-the-art generative edits—this technology helps combat deepfake misinformation while safeguarding creator rights. The proposed cross-attention architecture and psychoacoustic loss represent technical innovations that balance robustness with perceptual quality, offering practical tools for content provenance tracking and copyright enforcement in real-world applications.
Potential dual-use implications necessitate ethical consideration: while designed to enhance media integrity, similar watermarking systems could theoretically enable covert tracking or censorship if misapplied. We advocate for transparent deployment practices and policy frameworks to prevent abuse, emphasizing watermarking’s role in supporting—not undermining—digital rights. Researchers should consider these societal impacts when developing attribution technologies, particularly regarding user consent and detection transparency.
@@ Appendix A Implementation Details
Model Architecture. Following AudioSeal (San Roman et al., 2024), we leverage a pair of convolutional encoder-decoder models that operate on the waveform domain to generate and detect watermarks. The watermark generator consists of a waveform encoder and decoder, both utilizing components from EnCodec Defossez et al. (2022). The encoder employs a 1D convolution containing 32 channels and kernel size 7, then uses four convolutional blocks. Each of these blocks contains a residual unit and a down-sampling layer, utilizing convolution with stride $S$ and kernel size $K=2S$. The residual unit contains two kernel-3 convolutions with a skip-connection, with channels doubling during down-sampling. The encoder ends with a two-layer LSTM and final 1D convolution having kernel size 7 and 128 channels. The stride $S$ values used are (2, 4, 5, 8), and residual units utilize the Exponential Linear Unit (ELU) as nonlinear activation. The decoder reflects the encoder’s structure but employs transposed convolutions, with strides in opposite order.
The detector contains an encoder, transposed convolution, and linear layer for the detection head. The encoder utilizes the same structure as the one from the generator but with different weights. The transposed convolution includes $H$ output channels and upscales the activation map to match the original audio dimension, yielding an activation map shaped $(T,H)$. This frame-level latent state serves both the detection head and the message decoding head. For the detection head, the linear layer serves to reduce the dimensionality of $H$ to two, followed by a softmax function that produces per-sample detection probability scores. Additional details on the message decoding head are provided in the next section.
@@
Perceptual Loss. Following AudioSeal (San Roman et al., 2024), we employ a pool of perceptual loss terms to guide the model to learn the watermarking task. The perceptual loss terms include: i) multi-scale mel-spectrogram loss $\mathcal{L}_{\text{msspec}}$ that computes both L1 loss on linear-scale mel-spectrograms and MSE loss on log-scale mel-spectrograms across multiple FFT window sizes (from $2^{6}$ to $2^{11}$), with each scale weighted by $\sqrt{2^{i}-1}$ for scale $i$; ii) feature matching loss $\mathcal{L}_{\text{feat}}$ that minimizes L1 distance between intermediate feature maps extracted from the adversarial discriminator, averaged across all layers. iii) waveform L1 loss $\mathcal{L}_{\text{waveform}}$ that minimizes L1 distance between the original waveform and the watermarked waveform; iv) our proposed psychoacoustic-aligned time-frequency masking loss $\mathcal{L}_{\text{mask}}$, which captures the weighted $\ell_{2}$ loss in the mel-spectrograms domain, using a per-tile weight obtained through simulation of the masker energy decaying in time-frequency domain.
@@
Evaluation Setup. For the attribution experiment, following San Roman et al. (2024), we set up a pool of $N$ potential users, where each user’s sequential bit message is derived from their ordinal number, i.e., the $i$-th user’s message is given by $w_{i}=\text{binary}(i)$. This creates a space of all potential messages $E_{W}$. The attribution is achieved by retrieving the message in $E_{W}$ with the smallest Hamming distance. Compared to bit-wise message accuracy, this metric provides a more practical measure of message retrieval performance, aligning better with real-world watermarking applications and attribution use cases. The message length (in bits) determines the maximum number of uniquely identifiable users. Specifically, for a message of $K$ bits, the system can support up to $2^{K}$ unique users. In our experiments, we use a 16-bit message, enabling attribution across $5362^{16}=65,\!$ 536 users at maximum.

### v3

@@ 5.1 Experimental Setup
Following prior works (San Roman et al., 2024; Chen et al., 2023), we use a sampling rate of 16 kHz and one-second mono samples for training ($T=16000$) under 16 diverse audio editing transformations. We train the models on a mixed audio dataset of 4100 hours containing speech (3016 hours VoxPopuli (Wang et al., 2021) and 100 hours LibriSpeech (Panayotov et al., 2015)), music (9 hours MusicCaps (Agostinelli et al., 2023) and 880 hours Free Music Archive (Defferrard et al., 2016)), and sound effects (98 hours AudioSet (Gemmeke et al., 2017)). For evaluation, we use a held-out test set from MusicCaps of size 100. For each audio file, we embed 100 distinct messages, resulting in 10k unique watermarked audio samples. Each of these samples is then subjected to 16 different audio transformations, leading to a total of 160k evaluated instances. The audio duration is set as 5s by default in evaluation. The loss weights are set as: $\lambda_{\text{TF}}=1$, $\lambda_{\text{adv}}=1$, $\lambda_{\ell_{1}}=0.1$, $\lambda_{\text{msspec}}=2$, $\lambda_{\text{detect}}=\lambda_{\text{message}}=10$. We use the Adam optimizer (Kingma and Ba, 2015) with learning rate 1e-5, $\beta_{1}=0.4$, $\beta_{2}=0.9$, and Exponential Moving Average (EMA) (Tarvainen and Valpola, 2017) with decay factor of $0.99$ updated at every step. We train for 73k steps with batch size 16 and latent size $H=32$. To boost the sampling of the transformation efficiency, we update the sampling probability of each transformation every 1000 steps on the validation set, adjusting it based on the validation accuracy of each transformation.
For the attribution experiment, we follow the simulation protocol in San Roman et al. (2024) that defines a message pool of size NN ($\in\{100,1000,10000\}$), where each message is uniquely associated with a different user. As $N$ increases, the message length (in bits) also grows, making attribution more challenging due to the increased complexity of distinguishing individual messages. During decoding, the message is retrieved from the pool by selecting the one with the closest Hamming distance. We define attribution accuracy as the fraction of correctly attributed messages among all audio inputs that were positively detected as watermarked. We compare to four state-of-the-art baselines: AudiowMark (Westerfeld, 2020), WavMark (Chen et al., 2023), TimbreWM (Liu et al., 2023a), AudioSeal (San Roman et al., 2024). Please refer to App. A for more details.
@@ 5.2 Detection and Attribution Effectiveness
Robustness to Standard Edits. We present the detection and attribution performance across a comprehensive suite of standard audio editing operations in Table 1. For detection, we report both the overall accuracy and the true/false positive rates (TPR/FPR). For attribution, we report the averaged performance across different user numbers (see Figure 6 for the decomposed result). Our method achieves new state-of-the-art performance on both tasks, maintaining high detection accuracy (99.19% average) and attribution accuracy (93%) on average across transformations. Notably, while AudioSeal achieves strong detection performance (97.1% average), it struggles to perform effective attribution (39% average). The traditional approach, AudiowMark, exhibits more balanced detection-attribution trade-offs (around 88% attribution) but lower overall detection performance (92.9% average). Under the speed-change transformation, our method achieves 99.5% detection accuracy while all other methods degrade to near random-guess levels (50-61%). While attribution remains challenging under speed change, we show that the attribution performance can be partially restored via a speed-reversion layer (see App. C.3.1 for details). Figure 6 further validates that XAttnMark consistently maintains high attribution accuracy even with increasing message pool sizes. Furthermore, we conduct statistical tests to validate the statistical significance of the improvement in attribution across different transformations (App. C.14 and App. C.3 for more details). In summary, we demonstrate that XAttnMark is more robust under standard edits for both detection and attribution.
@@
Robustness to Generative Edits. Beyond the standard audio transformations that are seen during training, audio generative editing is one particularly challenging transformation that watermarking systems might have to endure at deployment (Liu et al., 2024a). To simulate this, we use two state-of-the-art audio generative models, AudioLDM2 (Liu et al., 2024a) and Stable Audio (Evans et al., 2024), with a text-guided DDIM inversion method proposed in ZETA (Manor and Michaeli, 2024). We test various editing strengths $t\in\{10,70,110\}$, which represent the diffusion forward step when using DDIM inversion. As shown in Table 2, AudiowMark and WavMark degrade to a random-guess level performance, while TimbreWM and AudioSeal show inferior performance across different editing strengths, around 50-60%. In contrast, our method maintains consistently high detection accuracy, averaging 91-94% across all editing strengths, and is consistent across both generative models. While XAttnMark demonstrates strong detection robustness (e.g., >>90% accuracy) against generative editing attacks in a zero-shot manner, achieving similarly robust attribution performance under such strong edits remains a challenge and an important direction for future work. Nevertheless, our detection performance marks a significant advancement. This demonstrates that XAttnMark generalizes better to unseen generative edits compared to existing methods. Please refer to the App. C.4 for more details.
Table 3: Detection performance and audio quality metrics after HSJA-based (black-box) adversarial attacks on the waveform and spectrogram domains on AudioMarkBench dataset (Liu et al., 2024b).
@@
Figure 3: Watermark-removal attacks under three different knowledge settings. PESQ is measured between attacked audios and genuine ones (PESQ << 4 indicates strong quality degradation).
Robustness to Adversarial Watermark Removal. Recent work by Liu et al. (2024b) demonstrates that adversarial perturbations can be crafted to make watermarks undetectable, even with only black-box access to the detection model. To evaluate robustness against such attacks, we employ the black-box HopSkipJumpAttack (HSJA) (Chen et al., 2020), which iteratively crafts minimal perturbations in either the waveform or spectrogram domain to fool the detector. As shown in Table 3, we first evaluate HSJA attacks in both domains with a query budget of $Q=100$. While the attacks successfully reduce detection accuracy, the attacks also significantly degrade audio quality in the process. In the waveform domain, increasing the query budget from $Q=100$ to $Q=500$ further reduces detection accuracy from 0.68 to 0.47. However, across all experimental configurations, XAttnMark consistently outperforms AudioSeal in terms of both detection robustness and perceptual quality preservation. Aligning with San Roman et al. (2024), we further evaluate XAttnMark against white-box, semi-black-box (using a re-trained surrogate model with the same architecture but different initialization), and Gaussian noise attacks, with results shown in Figure 3. The results reveal that model vulnerability increases with attacker knowledge, highlighting the importance of keeping detector models proprietary to mitigate watermark removal threats.
Extending to Watermark Localization. We further extend XAttnMark to perform watermark localization following the Brute Force Detection (BFD) method used in WavMark (Chen et al., 2023). This is achieved by employing a sliding window detection mechanism. Given that our model operates on 1-second audio segments and incorporates transformations robust to time shifts, we can distribute the per-segment detection probability to a finer, per-frame level by utilizing multiple overlapping detection windows. As shown in Figure 4, XAttnMark achieves localization performance comparable to AudioSeal and significantly outperforms WavMark, in terms of both sample-level F1 score and Intersection Over Union (IoU) metrics. Furthermore, we also demonstrate the effectiveness in the case where the duration of the watermark segment is fixed at 1 second, and the total audio length varies from 2 seconds to 10 seconds. Please refer to the App. C.10 for more results.
Figure 4: Watermark localization results across different durations of watermarked windows in terms of sample-level F1 score and Intersection Over Union (IoU) metrics ($\uparrow$ is better).
Table 4: Audio Quality Metrics. Comparison of objective perceptual quality metrics across watermarking methods.
Methods SI-SNR $\uparrow$ PESQ $\uparrow$ STOI $\uparrow$ ViSQOL $\uparrow$ L($\bm{\delta_{w}}$) $\downarrow$
@@
5.3 Quality and Stealthiness Assessment
We evaluate the following objective quality metrics of the watermarked audio: Scale Invariant Signal to Noise Ratio (SI-SNR), as well as PESQ (Rix et al., 2001), ViSQOL (Hines et al., 2012) and STOI (Taal et al., 2010). Furthermore, we also report the loudness of watermark residual $L(\bm{\delta_{w}})$ according to ITU-R BS.1770-4 (International Telecommunication Union, 2011), which is an important metric for watermarking stealthiness, especially under residual-based detection (Yang et al., 2024). As shown in Table 4, XAttnMark achieves competitive performance across all perceptual quality metrics. While WavMark achieves the highest SI-SNR (36.35 dB), XAttnMark excels in perceptual quality measures, achieving the best STOI score (1.000) and the lowest watermark residual loudness (-54.63 LUFS). For the PESQ and ViSQOL scores, XAttnMark achieves competitive performance, with a PESQ score of 4.43 and a ViSQOL score of 4.56. We also conducted a subjective listening test using the MUSHRA protocol (International Telecommunication Union, 2014) with 12 participants, and the results suggest that XAttnMark achieves a comparable quality with AudioSeal scoring around 91 (while ground truth scores around 95, more details in the App. C.1 and App. C.9). In summary, XAttnMark achieves superior utility while maintaining competitive perceptual quality.
@@ 5.4 Ablation Study
We first conduct ablation on the cross-attention module and temporal modulation layer to showcase their effectiveness for boosting learning in message decoding. As shown in Figure 6, our approach demonstrates superior learning efficiency, reaching approximately 98% accuracy. When removing the modulation component, performance drops significantly to around 60%. More dramatically, ablating the cross-attention mechanism causes accuracy to drop to random‑guess levels (50%), highlighting its crucial role in the architecture. Our ablation studies highlight the contributions of key architectural components. Notably, the partial parameter sharing enabled by the cross-attention mechanism is particularly impactful for improving message decoding efficiency and accuracy. While the temporal conditioning module also contributes, the ability of the detector to directly attend to shared message embeddings via cross-attention appears to be a more dominant factor in achieving robust attribution compared to ablations without it. Furthermore, we ablate the proposed TF loss $\mathcal{L}_{\text{TF}}$ and the adaptive bandwidth with constant weight $\gamma=1$. As shown in Table 4, both of them contribute to perceptual quality.
@@ 6 Conclusion
In this paper, we propose a neural audio watermarking framework that achieves both reliable watermark detection and accurate message attribution under challenging transformations, including generative editing. By integrating partial parameter sharing, a cross-attention-based detector, and a temporal conditioning module, our approach improves robustness and decoding accuracy compared to existing neural watermarking methods. Additionally, the psychoacoustic-aligned TF masking loss enables quality-preserving watermarking. To our knowledge, this is the first method to demonstrate non-trivial watermark detection under generative editing, marking a significant step toward reliable content attribution in generative AI settings. Future work includes developing stronger watermarking for robust attribution under speeding and various generative editing, and leveraging spatial psychoacoustics to improve the perceptual quality and support spatial audio watermarking.
Acknowledgments
This work was supported in part by Dolby Laboratories Inc. and Lehigh University. In this work, Yixin Liu and Lichao Sun were partially supported by the National Science Foundation Grants CRII-2246067, ATD-2427915, NSF POSE-2346158, NSF POSE-2449280, and Lehigh Grant FRGS00011497. This work used the Delta and DeltaAI systems at the National Center for Supercomputing Applications (NCSA) and the Bridges-2 system at the Pittsburgh Supercomputing Center (PSC) through allocation CIS240308 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program (Boerner et al., 2023), which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. We also thank Leslie Famularo from Dolby Laboratories Inc. for helpful discussions and the anonymous reviewers for their constructive feedback.
@@ Acknowledgments
This research addresses pressing concerns regarding media authenticity and intellectual property protection stemming from the rise of generative AI. Our work introduces an improved audio watermarking system. The system’s value lies in its ability to reliably detect embedded identifying marks (watermarks) and accurately trace their origin, even when audio has been substantially altered by sophisticated AI tools. This capability makes our technology a valuable asset in combating deepfake audio and the spread of misinformation. Furthermore, it offers enhanced protection for the rights of audio content creators.
Like many technologies, advanced watermarking holds the potential for dual use. Although designed to protect and authenticate, such systems could potentially be misused by malicious actors or institutions for purposes such as unauthorized tracking or censorship if implemented irresponsibly. Recognizing these risks, we strongly advocate for transparency and informed user consent in the deployment of such systems. The development of clear ethical guidelines and robust policies is crucial to prevent abuse.
@@ Appendix A Implementation Details
Model Architecture. Following AudioSeal (San Roman et al., 2024), we leverage a pair of convolutional encoder-decoder models that operate on the waveform domain to generate and detect watermarks. The watermark generator consists of a waveform encoder and decoder, both utilizing components from EnCodec (Défossez et al., 2023). The encoder employs a 1D convolution containing 32 channels and a kernel size of 7, then uses four convolutional blocks. Each of these blocks contains a residual unit and a down-sampling layer, utilizing convolution with stride $S$ and kernel size $K=2S$. The residual unit contains two kernel-3 convolutions with a skip-connection, with channels doubling during down-sampling. The encoder ends with a two-layer LSTM and a final 1D convolution having a kernel size of 7 and 128 channels. The stride $S$ values used are (2, 4, 5, 8), and residual units utilize the Exponential Linear Unit (ELU) as a nonlinear activation. The decoder reflects the encoder’s structure but employs transposed convolutions, with strides in the opposite order.
The detector contains an encoder, transposed convolution, and linear layer for the detection head. The encoder utilizes the same structure as the one from the generator, but with different weights. The transposed convolution includes $H$ output channels and upscales the activation map to match the original audio dimension, yielding an activation map shaped $(T,H)$. This frame-level latent state serves both the detection head and the message decoding head. For the detection head, the linear layer serves to reduce the dimensionality of $H$ to two, followed by a softmax function that produces per-sample detection probability scores. To obtain the sample-level detection probability from frame-level logits, we apply temporal averaging (mean pooling) across the time dimension. Additional details on the message decoding head are provided in the next section.
@@
Perceptual Loss. Following AudioSeal (San Roman et al., 2024), we employ a pool of perceptual loss terms to guide the model to learn the watermarking task. The perceptual loss terms include: i) multi-scale mel-spectrogram loss $\mathcal{L}_{\text{msspec}}$ that computes both L1 loss on linear-scale mel-spectrograms and MSE loss on log-scale mel-spectrograms across multiple FFT window sizes (from $2^{6}$ to $2^{11}$), with each scale weighted by $\sqrt{2^{i}-1}$ for scale $i$; ii) feature matching loss $\mathcal{L}_{\text{feat}}$ that minimizes L1 distance between intermediate feature maps extracted from the adversarial discriminator, averaged across all layers. iii) waveform L1 loss $\mathcal{L}_{\text{waveform}}$ that minimizes L1 distance between the original waveform and the watermarked waveform; iv) our proposed psychoacoustic-aligned TF masking loss $\mathcal{L}_{\text{mask}}$, which captures the weighted $\ell_{2}$ loss in the mel-spectrograms domain, using a per-tile weight obtained through simulation of the masker energy decaying in the time-frequency domain.
@@
Evaluation Setup. For the attribution experiment, following San Roman et al. (2024), we set up a pool of $N$ potential users, where each user’s sequential bit message is derived from their ordinal number, i.e., the $i$-th user’s message is given by $w_{i}=\text{binary}(i)$. This creates a space of all potential messages $E_{W}$. The attribution is achieved by retrieving the message in $E_{W}$ with the smallest Hamming distance. Compared to bit-wise message accuracy, this metric provides a more practical measure of message retrieval performance, aligning better with real-world watermarking applications and attribution use cases. The message length (in bits) determines the maximum number of uniquely identifiable users. Specifically, for a message of $K$ bits, the system can support up to $2^{K}$ unique users. In our experiments, we use a 16-bit message, enabling attribution across $2^{16}=65,\!536$ users at maximum.